OpenStudy (anonymous):

I WILL MEDAL ANYONE PLEASE PLEASE HELP ME Which of the following represents 6 x to the 3 fourths power in radical form? cubed root of 6 x to the fourth power fourth root of 6 x cubed 6 cubed root of x to the fourth power 6 fourth root of cubed

OpenStudy (studygurl14):

like this \[6x^3\] or this \[(6x)^3\]

OpenStudy (anonymous):

6x^3/4

OpenStudy (anonymous):

isn't the answer 6 fourth root to x cubed?

OpenStudy (studygurl14):

oops, sorry forgot the 4. But anyway, is does the exponent apply to the 6 too, or just the x?

OpenStudy (anonymous):

im confused??

OpenStudy (anonymous):

I figured this one out... can you help me with another one Rationalize the denominator of 7 over quantity of 3 plus square root of 3

OpenStudy (anonymous):

the possible answers are 7 square root of 3 over 12 quantity of 21 minus 3 over 6 7 square root of 3 over quantity of 3 square root of 3 end root minus 3 quantity of 21 minus 7 square root of 3 over 6

OpenStudy (anonymous):

it looks like 7/ 3 + sqrt3

OpenStudy (studygurl14):

OH OKAY HOLD ON

OpenStudy (studygurl14):

oops cap lock was on

OpenStudy (anonymous):

haha its okay!

OpenStudy (studygurl14):

sorry it took so long...computer having trouble..hold on one more sec

OpenStudy (studygurl14):

\[\frac{7}{3+\sqrt{3}}\] \[\frac{7}{3+\sqrt{3}}\times \frac{3-\sqrt{3}}{3-\sqrt{3}}\] \[\frac{21-7\sqrt{3}}{9-\sqrt{9}}=\frac{21-7\sqrt{3}}{9-3}\] \[=\frac{21-7\sqrt{3}}{-6}\] So..what's the answer?

OpenStudy (studygurl14):

sorry ignore the negative sign in front of the six